OpenStudy (anonymous):

3. Find the first four terms of the sequence represented by the expression. 3n + 5 3, 8, 14, 20 8, 11, 14, 17 3, 6, 9, 12 0, 8, 11, 14 I really just need help understanding this... none of the answers look right.

OpenStudy (anonymous):

@Ilovepurple42043

OpenStudy (anonymous):

put n=1,2,3,4,5

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

first term=3*1+5=3+5=8 second term=3*2+5=6+5=11 and so on

OpenStudy (anonymous):

O, I get it; thanks so much!

OpenStudy (anonymous):

yw